OpenStudy (anonymous):

Find the slope of the line tangent to the curve at t=(pi/6). Where x=4cost, y=4sint 0

OpenStudy (anonymous):

Hint:\[\frac{dy}{dx}=\frac{\dfrac{dy}{dt}}{\dfrac{dx}{dt}}=\frac{\dfrac{d}{dt}[4\sin t]}{\dfrac{d}{dt}[4\cos t]}=\cdots\]